### v1

@@ Abstract
Motivated by this complexity, a mixture of expert LLMs tailored to specific financial domains could offer a more comprehensive understanding for intricate financial tasks. In this paper, we present the FinTeamExperts, a role-specialized LLM framework structured as a Mixture of Experts (MOEs) for financial analysis. The framework simulates a collaborative team setting by training each model to specialize in distinct roles: Macro Analysts, Micro analysts, and Quantitative Analysts. This role-specific specialization enhances the model’s ability to integrate their domain-specific expertise. We achieve this by training three 8-billion parameter models on different corpus, each dedicated to excelling in specific finance-related roles. We then instruct-tune FinTeamExperts on downstream tasks to align with practical financial tasks. The experimental results show that FinTeamExperts outperform all models of the same size and larger on three out of four datasets. On the fourth dataset, which presents a more complex task, FinTeamExperts still surpass all models of the same size. This highlights the success of our role-based specialization approach and the continued training approach for FinTeamExperts.
@@ 1 Introduction
Our methodology follows a two-phase training process. In the first phase, the models are pre-trained on a curated corpus of role-related data to establish a strong foundational understanding of financial concepts. In the second phase, the models undergo specific instruct-tuning, where they are fine-tuned to generate outputs that align closely with practical downstream tasks in the finance domain. The supervised instruct-tuning ensures that FinTeamExperts are skilled at applying knowledge to different real-world scenarios, making them effective for complex financial analysis and decision-making tasks.
@@ 2.1 Framework Formulation
The core of the MOE is the routing mechanism of inputs to Experts. We set up a group of experts, corresponding to macro experts, micro experts, and quant experts. A hard gating mechanism is used to select one expert from each group, and their outputs are then weighted via another soft gating mechanism. The routing gate can be expressed as $g(x_{i})$ representing a probability distribution over the experts.
@@
We define $y_{i}$ as the output for the input token $x_{i}$, which is a weighted sum of the selected experts’ outputs. $E_{j}(x_{i})$ is the output of expert $j$ for input $x_{i}$, where each expert processes the input differently. $K$ is the total number of experts available, for instance, three in our study. Typically, only a subset of these experts is used for each input but we have it as a weighted contribution.
We define the total loss for the MOE model as $\mathcal{L}_{\text{task}}$, including both the task-specific loss and the regularization term. The task-specific loss (e.g., cross-entropy) is used for training the model on the primary task.
@@
where $\lambda$ is a regularization coefficient controlling the impact of the entropy term. $H(g(x))$ is the entropy of the gate function output, used to encourage balanced utilization of experts.
@@ 3.2 Downstream Datasets
User What is the sentiment of the following financial post: Positive, Negative, or Neutral? Text: Whats up with \$LULU? Numbers looked good, not great, but good. I think conference call will instill confidence.
Assistant The sentiment is neutral
User Examine the data and tweets to deduce if the closing price of \$cvx will boost or lower at 2017-02-21. Kindly confirm either Rise or Fall. Context: date,open,high,low,close,adj-close,inc-5,inc-10,inc-15,inc-20,inc-25,inc-30 2017-02-06,0.3,0.5,-0.5,-0.5,-0.5,-0.7,0.6,1.2,1.6,2.1,2.5 2017-02-07,1.4,1.5,-0.3,-1.4,-1.4,0.8,1.6,2.4,2.8,3.3,3.7 2017-02-08,-0.4,0.2,-1.1,0.2,0.2,0.7,0.9,2.0,2.5,2.9,3.3 2017-02-09,-0.6,0.3,-0.6,0.6,0.6,0.1,-0.1,1.1,1.7,2.1,2.5 2017-02-10,-0.3,0.4,-0.6,0.7,0.7,-0.7,-0.8,0.3,0.9,1.2,1.7 2017-02-13,-0.5,0.1,-0.8,0.8,0.8,-1.3,-1.4,-0.6,-0.0,0.3,0.8 2017-02-14,0.1,0.5,-0.7,-1.2,-0.3,-0.6,-0.9,-0.4,0.1,0.5,1.0 2017-02-15,-0.3,0.6,-0.3,0.0,0.0,-0.3,-0.7,-0.7,-0.0,0.4,0.8 2017-02-16,2.1,2.1,-0.2,-1.7,-1.7,1.3,0.9,0.7,1.5,2.0,2.3 2017-02-17,0.1,0.2,-1.0,-0.3,-0.3,1.3,1.0,0.9,1.7,2.2,2.5 2017-02-06: #dividend growth investing at #work - the streak continues! #investing #dividends \$cvx |rt AT_USER chevron corporation \$cvx shares bought by los angeles capital management & equity research inc. 2017-02-07: rt AT_USER 4 energy stocks that are ticking time bombs \$cvx \$oxy \$cop |rt AT_USER mike liss of AT_USER century (\$twvlx) talked about \$cvx \$apc \$mdlz \$mdt & \$mrk during "hold it or fold it" in tod 2017-02-08: rt AT_USER dynamic capital management ltd reduces stake in chevron corporation \$cvx |dynamic capital management ltd reduces stake in chevron corporation \$cvx |rt AT_USER #chevron partners zinox for co 2017-02-09: Answer:
Assistant The prediction is Rise
@@ 4.1 Financial LLMs
A survey regarding FinLLMs [16] shows that research in this area remains limited.
FinBERT [5] adapts BERT for financial sentiment analysis in two steps. First, it is pre-trained on financial documents using a subset of Reuters’ TRC2 dataset, enhancing its understanding of financial terminology. Second, a dense layer is added to the last hidden state of the classification token (CLS) and fine-tuned with the Financial PhraseBank (FPB) dataset. FinBERT achieves outstanding results in financial sentiment analysis, surpassing state-of-the-art models, but is limited to this task and performs poorly on others.

### v2

@@ Abstract
Motivated by this complexity, a mixture of expert LLMs tailored to specific financial domains could offer a more comprehensive understanding for intricate financial tasks. In this paper, we present the FinTeamExperts, a role-specialized LLM framework structured as a Mixture of Experts (MOEs) for financial analysis. The framework simulates a collaborative team setting by training each model to specialize in distinct roles: Macro Analysts, Micro Analysts, and Quantitative Analysts. This role-specific specialization enhances the model’s ability to integrate their domain-specific expertise. We achieve this by training three 8-billion parameter models on different corpus, each dedicated to excelling in specific finance-related roles. We then instruct-tune FinTeamExperts on downstream tasks to align with practical financial tasks. The experimental results show that FinTeamExperts outperform all models of the same size and larger on two out of four datasets. On the stock prediction, which presents a more complex task, FinTeamExperts still surpass all models of the same size. This highlights the success of our role-based specialization approach and the continued training approach for FinTeamExperts.
@@ 1 Introduction
Our methodology follows a two-phase training process, as shonw in the Figure 1. In the first phase, the models are pre-trained on a curated corpus of role-related data to establish a strong foundational understanding of financial concepts. In the second phase, models are combined using routing gates and undergo instruct-tuning. This fine-tuning process helps the models produce outputs that are well-aligned with real-world tasks in the finance domain. The supervised instruct-tuning ensures that FinTeamExperts are skilled at applying knowledge from different experts to different real-world scenarios, making them effective for complex financial analysis and decision-making tasks.
@@ 2.1 Framework Formulation
The core of the MOE is the routing mechanism of inputs to Experts. We set up a group of experts, corresponding to macro experts, micro experts, and quant experts. A hard gating mechanism is used to select one expert from each group, and their outputs are then weighted via another soft gating mechanism. The routing gate can be expressed as $g(x_{i})g(x_{i})=\big{(}g_{1}(x_{i}),g_{2}(x_{i}),\dots,g_{K}(x_{i})\big{)}$ representing a probability distribution over the experts.
@@
In a simplified three-role gating framework, we can denote the output $y_{i}$ for an input token $x_{i}$, as a weighted sum of the outputs from the selected experts. Each expert $E_{j}(x_{i})$ of expert $j$ produces a unique output for teh input $x_{i}$. $K$ is the total number of experts available, for instance, three in our study. Typically, only a subset of these experts is used for each input but we have it as a weighted contribution.
To build a adaptable framework where a flexible number (e.g., K) of experts caon contribute to each expert category. The weighted output $y_{i}$ is thus adjusted as a sum over all roles and their respective experts, with each category weighted by a corresponding gating functions. This is defined as:
$y_{i}=\sum_{r\in\{\text{Macro},\text{Micro},\text{Quant}\}}\sum_{j=1}^{K}g_{r}(x_{i})\cdot h_{r,j}(x_{i})\cdot E_{r,j}(x_{i})$ (3)
where $r$ denotes the role (e.g., Macro, Micro, Quant), and $g_{r,j}(x_{i})$ is the gating function that assigns weights to each expert. In particular, we set
$h_{r,j}=\text{GumbelSoftmax}(W_{g}\cdot x_{i}),$ (4)
for all $r$ to select one expert from each group of experts.
For inference, we define the inference head $\hat{y}_{t}=C(y_{j<t})$, which processes outputs from previous tokens to generate the next prediction.
The overall loss for the FinTeamMOE model, $\mathcal{L}_{\text{task}}$, includes both the task-specific loss and a regularization term. The task-specific loss, such as cross-entropy, is used for training the model on the primary task:
@@
where $\lambda$ is a regularization coefficient controlling the influence of the entropy term $H(g(x))$. This entropy term promotes balanced utilization of experts by encouraging the gate function to distribute weights across experts effectively.0
@@ 3.2 Downstream Datasets
User What is the sentiment of the following financial post: Positive, Negative, or Neutral?
Text: What’s up with \$LULU? Numbers looked good, not great, but good. I think conference call will instill confidence.
Assistant The sentiment is neutral.
User Examine the data and tweets to deduce if the closing price of \$cvx will boost or lower at 2017-02-21. Kindly confirm either Rise or Fall.
Context: date,open,high,low,close,adj-close,inc-5,inc-10,inc-15,inc-20,inc-25,inc-30
2017-02-06,0.3,0.5,-0.5,-0.5,-0.5,-0.7,0.6,1.2,1.6,2.1,2.5
2017-02-07,1.4,1.5,-0.3,-1.4,-1.4,0.8,1.6,2.4,2.8,3.3,3.7
2017-02-08,-0.4,0.2,-1.1,0.2,0.2,0.7,0.9,2.0,2.5,2.9,3.3
2017-02-06: #dividend growth investing at #work - the streak continues! #investing #dividends \$cvx
rt AT_USER chevron corporation \$cvx shares bought by los angeles capital management & equity research inc.
2017-02-07: rt AT_USER 4 energy stocks that are ticking time bombs \$cvx \$oxy \$cop
rt AT_USER mike liss of AT_USER century (\$twvlx) talked about \$cvx \$apc \$mdlz \$mdt & \$mrk during "hold it or fold it"
2017-02-08: rt AT_USER dynamic capital management ltd reduces stake in chevron corporation \$cvx
dynamic capital management ltd reduces stake in chevron corporation \$cvx
rt AT_USER #chevron partners zinox for co
2017-02-09: Answer:
Assistant The prediction is Rise.
@@ 3.5 Ablation Study
We conduct a perplexity analysis on three role-oriented training models and our FinTeamExperts model, using mini-test samples to assess their perplexity scores. Figure LABEL:fig illustrates these results, with a benchmark perplexity score of approximately 15 for the vanilla Llama-3-8B model before it was trained with role-specific adaptations. Perplexity trends are plotted for the macro, micro, and quant models, while downstream learning results are shown for FinTeamExperts. As training progressed, perplexity scores for all three role-oriented models decreased and eventually converged around 6.5. Among these, the macro-role model achieved the lowest perplexity, with the quant-role model slightly higher, though the differences were minor.
FinTeamExperts, routed and fine-tuned for downstream financial tasks, began with an initial perplexity below 8 and converged to a value under 6. This indicates that FinTeamExperts is better adapted to the financial dataset, demonstrating enhanced understanding and alignment with the financial tasks.
@@ 4.1 Financial LLMs
Prior to the rise of LLMs, deep learning played a pivotal role in the financial domain, particularly in tasks like portfolio management [20] and risk management [21]. In portfolio management, deep learning can be utilized to enhance the optimization process by directly maximizing metrics such as the Sharpe ratio. This approach simplifies asset selection by effectively capturing correlations across different asset classes, thereby streamlining decision-making in constructing a balanced portfolio.
In the era of advanced LLMs, research specifically targeting financial language models (FinLLMs) remains relatively limited, as highlighted by a recent survey [22]. One notable example is FinBERT [5], which adapts the general-purpose BERT model for financial sentiment analysis using a two-step approach. First, it is pre-trained on financial texts, utilizing a subset of the Reuters TRC2 dataset to enhance its understanding of financial terminology. Next, a dense layer is added to the final hidden state of the classification token (CLS), and the model is fine-tuned using the Financial PhraseBank (FPB) dataset. FinBERT achieves outstanding performance in financial sentiment analysis, surpassing state-of-the-art models in this specific task. However, it remains limited in scope and does not perform as effectively on other financial tasks.
FinGPT [23] provides an end-to-end framework for training and applying financial large language models in the finance industry. It uses the efficient Low-rank Adaptation (LoRA) [24] technique to fine-tune open-source models like LLaMA and ChatGLM with around 50,000 samples. However, its evaluation is currently limited to finance classification tasks only.